OpenStudy (anonymous):

Which equation is correct? sec x° = opposite ÷ adjacent cot x° = opposite ÷ adjacent cosec x° = opposite ÷ adjacent sec x° = hypotenuse ÷ adjacent

OpenStudy (faiqraees):

\[\large\rm cotx = \frac{opposite}{adjacent} \]

OpenStudy (unklerhaukus):

sin = opp/hypo cos = adj/hypo tan = sin/cos = opp/adj sec = 1/cos csc = 1/sin cot = 1/tan

OpenStudy (anonymous):

thanks guys